if y is proportional to 2x-7,and y=-9 when x=2,find x when y=-15.Please show me step by step.
Answer by flame8855(424)

since y is proportional to 2x-7 then y= k(2x-7)
when x = 2 , y= -9
-9=k(4-7)
k= 3
so y=3(2x-7)
when y= -15
-15=3(2x-7)
2x-7=-5
x=1